OpenStudy (anonymous):

can anyone help me simplify the sqrt of -121 using complex numbers or something

OpenStudy (anonymous):

sure since \(\sqrt{121}=11\) you have \[\sqrt{-121}=\sqrt{-1\times 121}=\sqrt{-1}\sqrt{121}=11i\][

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw